rxcklesskaisher:

The new club sticker is in the shape of a triangle. It has a height of 16 inches and a base of 14 inches. The area of the club sticker is ____ square inches.

rxcklesskaisher:

1 attachment
AZ:

Area of a triangle = 1/2 * base * height

AZ:

The base is 14 The height is 16 Can you put that into the formula and find the area?

AZ:

Area = 1/2 * 14 * 16 Area = ???

rxcklesskaisher:

idk

AZ:

can you multiply the numbers?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Area = 1/2 * 14 * 16 Area = ??? \(\color{#0cbb34}{\text{End of Quote}}\) little confused written A = (1*14*16)/2 = ?

rxcklesskaisher:

it will be way 2 high

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Area = 1/2 * 14 * 16 Area = ??? \(\color{#0cbb34}{\text{End of Quote}}\) little confused written A = (1*14*16)/2 = ? \(\color{#0cbb34}{\text{End of Quote}}\) AZ this is your work - sorry my interfere

rxcklesskaisher:

im lost

snowflake0531:

14*16/2 put that into the calculator

rxcklesskaisher:

112

AZ:

That's your answer

rxcklesskaisher:

thanks @AZ @snowflake0531 and @jhonyy9

AZ:

No problem